# Self-Admitted GenAI Usage in Open-Source Software

### v2

@@ 3.2.1 GenAI-Assisted Tasks
Overall, our analysis identified 32 distinct task categories in which developers use GenAI tools in their workflows. Table II presents these categories along with their definitions and usage frequencies. Unsurprisingly, excluding PR-related activities, generation tasks dominated the landscape, with code generation being particularly prominent (105 instances). Translation followed with 50 instances, while optimization and maintenance tasks accounted for 34 and 26 instances, respectively.
As mentioned above, we distinguish between developer tasks and GenAI tasks. While Table II lists the GenAI tasks, we also want to discuss human tasks related to GenAI tasks. For example, in one commit message that we analyzed (EIII) the developer acknowledged that the code was written a bit hasty on previous release due to trust in GitHub Copilot. The developer tasks described in the commit message was bug fixing, while the initial task that the GenAI tool supported was code generation.
@@
Recent research has shown that using AI-generated PR descriptions reduces review time and increases PR merge rates [7]. We found that developers reused generated PR descriptions as part of their commit messages. As mentioned above, this approach was very common in one particular project, which contributed 1008 of such mentions to our sample. To illustrate this particular use case, we include an excerpt below (EIII). Interestingly, the linked contribution guidelines (EIII) do not discuss GenAI usage.
@@
In the following, we discuss the most prevalent supported tasks besides generating PR descriptions. As expected, code generation was one of the most common GenAI-supported tasks that we observed. Some self-admitted GenAI usage for code generation were straightforward, such as the following statement that we found in the source code comment documenting a method written in C#: This function was written with Chat-GPT (EIII).
Beyond code generation, developers used GenAI tools to generate other software artifacts, including test data or documentation. Besides generation, GenAI tools were also used to automate code review, for example as part of GitHub Actions workflows (EIII): Recently, the team has added ChatGPT to GitHub Actions to perform GenAI-based code review. The specific job can be found in the chatgpt-cr.yml file.
After generation, translation emerged as the second most prevalent task in our analysis. Most mentions referred to translation between natural languages, one mention referred to translation between programming languages. An important use case was internationalization, helping developers overcome language barriers (EIII): Due to my limited proficiency in English, all English document translations are currently provided by ChatGPT, including this sentence. The one mention related code translation documented the translation of existing Python to R code (EIII): The following R code was generated using ChatGPT based on the Python code. However, the developer at the same time asked others to support them in improving the code: If anyone can refactor it to something more readable or more sensible code, please let me know or submit as a pull request.
Code optimization represented the third largest category. Developers not only acknowledged GenAI tools usage but sometimes even thanked the tools in their commit messages (EIII): Forgot tabs. Thanks, Copilot. In addition to code, GenAI tools were also used to improve UI elements (EIII): I asked chatGPT to help me brainstorm improvements to some of the labels and hint text based on the Apple Human Interface Guidelines. I then edited them as human to improve them further. Interestingly, also in this case the developer asked other team members to review the generated content: Review and let me know if you think any are worse or weird. This, together with the human corrective actions triggered by GenAI actions we observed, points to the importance of human oversight in GenAI-assisted software development.
@@ 3.2.2 Generated Content Types
Project metadata Commit messages GenAI tools target commit messages. 1,003
Source files Whole methods GenAI tools target source code files, ranging from entire functions within a file. 47
Blocks within one source code file GenAI tools target source code files, spanning multiple blocks within a single source code file. 45
One block within one source code file GenAI tools target source code files, spanning one block within a single source code file. 39
Blocks within multiple source code files GenAI tools target source code files, spanning multiple source code files. 21
Whole files GenAI tools target source code files, ranging across the entire file. 12
Whole classes GenAI tools target source code files, ranging across the entire class in the file. 12
Project assets Documentation files GenAI tools target documentation files, which include technical documents in software projects. 106
Configuration files GenAI tools target configuration files, which are crucial for defining the operational parameters and settings of software applications. 24
Resource files GenAI tools target resource files, which include assets like images, localization strings, and other binary data. 5
@@
Categories Codes Definition #
@@
Removal of Copilot comment Indicating the deletion of comments or pieces of code initially suggested by a GenAI tool that are no longer relevant or correct. 2
@@
Warning Issuing cautions about potential issues in the code, such as security vulnerabilities, deprecated methods, or unstable features. 10
@@
GenAI Limitations Blame Copilot Specifically attributing errors or suboptimal code to suggestions made by a GenAI tool 3
@@
Our analysis identified three main categories of AI-generated content in open-source software projects organizing ten distinct codes (see Table IV). Although, as mentioned before, commit messages related to Copilot PR activities dominated our dataset with over 1,000 mentions from a single repository, examining the remaining data revealed important patterns. Developers frequently use GenAI tools to modify source files (176 mentions). However, other file types such as documentation and configuration files were also targeted (135 mentions).
@@
For project assets other than source code, GenAI was used, for example, to generate changelogs (EIII): Note: This changelog is improved by OpenAI ChatGPT from my broken English input. Another use case we observed was adding comments explaining options in a configuration file (EIII): These strings were provided by GitHub Copilot. I checked the first few, and they were correct.
@@ 3.2.3 Purposes of GenAI Usage
Summary RQ1: For the 1,292 GenAI mentions we analyzed, developers mainly used GenAI tools for code generation, natural language translation, and code refactoring. Source code and documentation files were the dominant generation targets. Acknowledgment of GenAI usage was a common purpose, sometimes combined with warnings about possible negative implications. Another important purpose was regulation (see RQ2). Our analysis revealed patterns of corrective actions following code generation. Our findings show that GenAI tools are actively used in open-source software, and that developers are working on guiding their usage.
@@ 4 Existing Guidelines for GenAI Usage
One topic that emerged while answering RQ1 is that some open-source projects have specific policies and guidelines around GenAI usage. Therefore, as part of RQ2, we investigated how projects regulate or recommend the usage of GenAI tools. In addition to analyzing the policies and guidelines, we conducted a survey with open-source developers to understand their views on GenAI regulation.
@@ 4.1 Method
Using our sample of GenAI mentions, we found 28 mentions related to policies and usage guidelines around GenAI tool usage. We grouped them into two groups: (1) exclusion of usage within the project and (2) regulation of usage within the project. Table VI presents detailed examples drawn from 13 documentation files (e.g., CONTRIBUTING.md) and commit messages from 12 GitHub repositories, where the last column indicates the number of mentions identified in the software artifact.
First, we closely examined these policies and usage guidelines to understand how exactly projects regulate GenAI usage. Second, we conducted a developer survey that included excerpts from the policies and guidelines we found. The primary goals of the survey were to: (1) collect developer perceptions on the need for GenAI tool guidance (e.g., documenting prompts or annotating generated content) and understand the actions taken on this content before integration or publication; and (2) investigate the rationale behind policies and usage guidelines. To investigate the second part, we asked participants if they contribute to one of the repositories from which we extracted policies and guidelines (see Table VI) and then showed the corresponding guidelines, asking them to elaborate on the rationale behind them. In this way, we received feedback on PVI and GVI. For developers who did not identify as contributors to one of the repositories, we showed them PVI, PVI, and GVI, asking for their feedback on those guidelines. We share the complete questionnaire as part of our replication package.
Of the 12 GitHub repositories that contained policies, seven used the GitHub Discussions feature, allowing us to gather direct developer feedback. To also cover repositories without this feature and repositories for which we did not receive responses, we reached out to 30 project developers using contact details found outside of GitHub (e.g., personal websites or social media). We received eight survey responses, which we analyzed using a combination of open coding and card sorting. Informed consent was obtained.
@@ 4.2 Results
Table VI: Policies and guidelines for GenAI usage in software projects (RQ2): exclusion of usage within the project and regulation of usage within the project; the last column (#) shows the number of mentions in the corresponding documentation file or commit message.
Purpose ID Repository Excerpt #
excl P1 jqwik-team/
@@
G2 avaloniaui/
@@
G4 owasp/
wrongsecrets
G5 sitespeedio/
sitespeed.io
G6 spring-projects/
@@ 4.2.1 Policies and Usage Guidelines of GenAI Tools
As mentioned above, Table VI lists 13 software artifacts from 12 GitHub repositories that presented policies or usage guidelines for GenAI usage.
Policies: Policies PVI–PVI illustrate community decisions that exclude GenAI usage in the projects. Maintainers of the project jqwik-team/jqwik raised concerns related to the copyright situation around GenAI-generated content (PVI and PVI). Similarly, maintainers of shoelace-style/shoelace addressed ethical and licensing issues arising from the inclusion of GenAI-generated code (PVI). Regarding code reviews, maintainers of katsutedev/mal4j (PVI) and shred/acme4j (PVI) explicitly stated that contributions generated by GenAI are not acceptable. The project turms-im/turms (PVI) discouraged the use of GenAI-generated responses in discussions, citing concerns over the lack of critical thinking and responsibility. In addition, the maintainers proposed to incorporate indicators for identifying possible GenAI usage and suggested tool support, for example, a ChatGPT detector published on HuggingFace.111huggingface.co/Hello-SimpleAI/chatgpt-detector-roberta These regulations demonstrate how open-source communities are beginning to establish boundaries and safeguards to ensure responsible integration of GenAI tools within collaborative open-source software development environments.
Usage Guidelines: Guildelines GVI–GVI outline recommendations for the appropriate use of GenAI tools in software development workflows. For example, the maintainers of graycoreio/daffodil (GVI) require developers to disclose any use of GenAI as a prerequisite for submitting a pull request. The maintainers of avaloniaui/avalonia (GVI) encouraged the use of GenAI to help draft pull request descriptions summarizing the results of the code review. In hardisgroupcom/sfdx-hardis (GVI), maintainers recommended to use GenAI for Q&A support, particularly for troubleshooting deployment issues. The project spring-projects/spring-cli (GVI) promoted the use of GenAI to generate README.md files and has even developed GenAI-integrated tooling to support automated documentation rewriting. Meanwhile, maintainers of sitespeedio/sitespeed.io, theokanning/openai- java, and owasp/wrong-secrets (GVI, GVI, and GVI) advised caution when using GenAI, warning of potential inaccuracies and security risks, such as inadvertent secret leakage due to the fact that tool vendors use prompts for reinforcement learning. Overall, these usage guidelines reflect a growing awareness of both the opportunities and risks of GenAI tools in open-source software projects and the willingness of the maintainers to guide their usage.
@@ 4.2.2 Developer Survey on GenAI Governance
Based on the analysis of the before-mentioned policies and usage guidelines, we designed ten questions regarding (i) the necessity of GenAI tool guidance; (ii) the necessity of documenting prompts and their generated contents; (iii) actions on generated content before integrating; and (iv) rationale behind policies and usage guidelines of real-world GenAI tool. In the following, we will use $D$ to refer to individual developers that participated in our survey.
@@
Documenting Prompts and Generated Content: $D_{5,6,8}$ emphasize the importance of documenting prompts and generated content to ensure accountability in software projects. They suggest two methods to achieve this: (1) associating prompts with their functionality and sharing them under a CC BY 4.0 license, and (2) embedding prompts as code comments or in project documentation, supplemented by shared discussion platforms like ChatGPT’s shared links. Despite some developers considering prompt documentation unnecessary, the majority agreed that it is valuable to understand the extent of GenAI’s contributions to a project. This documentation is essential for assessing which code is potentially affected by copyright and licensing issues; it might also prove useful later maintenance activities.
@@
Actions on Generated Content Before Integration: Developers are, compared to manually written code, more likely to perform code reviews and license compliance checks on AI-generated content before integrating it into their projects. Three developers highlighted these practices as crucial steps to ensure the quality and compliance of GenAI-based contributions. Additionally, some developers indicated that they rely on automated tools, e.g. code quality checks or automated testing, to evaluate generated content. One developer noted the importance of adding comments to document generation context. Interestingly, $D_{2}$ explicitly stated that no additional actions are necessary, explaining that all content in the PR will be subjected to rigorous review and testing regardless. This response reflects the perspective that standard testing and code review are sufficient to ensure the quality of both AI-generated and manually created content.
Project-specific GenAI Usage Guideance: The feedback we received from open-source developers regarding GenAI tool guidance reflect a combination of ethical, legal, and practical considerations. For example, the project owner of jqwik-team/jqwik ($D_{3}$), described their decision to disallow the use of GenAI tools as an ethical decision due to all its collateral damages. This statement suggests a strong position against the potential implications of accepting AI-generated contributions, with a particular focus on copyright and ethics. The regulation in the accompanying contributor agreement (PVI) explicitly prohibits contributions created using GenAI tools, citing the unresolved legal implications of training AI models on public code repositories.
@@
Similarly, the project owner of owasp/wrongsecrets ($D_{2}$) focuses on the ethical risks of using GenAI when describing the rationale behind guideline GVI. They highlight the importance of vigilance when handling sensitive data, particularly in the wrongsecrets project. They reported: This is a recommendation meant for people using WrongSecrets, and it applies more broadly than WrongSecrets or even OWASP itself. You should be conscious about what data you share, and be vigilant that you don’t input sensitive data, since tenant boundaries are mirky at best. This calls a broader concern about how user input may be stored or reused by GenAI systems. The associated recommendation emphasizes that GenAI tools often rely on reinforcement learning, which could expose sensitive data to unintended parties.
@@
Guideline GVI, which requires the disclosure of AI usage in pull requests, received support from three developers ($D_{4}$, $D_{5}$, $D_{6}$). $D_{4}$ emphasized that disclosure depends on whether the backbone/key idea was generated by AI, while $D_{5}$ highlighted the importance of transparency to maintain license compliance. $D_{6}$ added that disclosing the percentage of GenAI involvement in contributions could reduce the likelihood of noise PRs by GenAI and improve code review efficiency. This reflects a growing recognition of the need for transparency in collaborative software development, where understanding the role of AI in contributions can improve accountability and ensure compliance.
@@
Opinions diverge considerably regarding policy PVI, which prohibits AI-generated code. $D_{4}$ opposed such restrictions, viewing them as unnecessary limitations that could stifle productivity and innovation. $D_{6}$ criticized the policy as being overly cautious, suggesting that asking contributors to disclose percentage of generated content is sufficient. $D_{7}$ supported the regulation, noting its alignment with their own concerns about the ethical and legal implications of using GenAI tools. $D_{5}$, pointed to ethical and legal ambiguities related to AI-generated code, describing them as maintainer’s main concerns. They specifically highlighted that AI models are likely trained on large datasets that include open-source codebases with various licensing terms.
$D_{4}$ and $D_{6}$’s feedback on policy PVI, which regulates the use of AI in community discussions, emphasizes concerns about the high false positive rates of AI identification tools and warns against deferring critical decisions to automation. $D_{5}$ argued that while LLMs are suitable for repetitive tasks and generic translations, they lack the creativity needed for meaningful contributions. This aligns with the cautionary tone of the regulation, which warns about overreliance on AI-generated content.

### v3

@@ 3.2.1 GenAI-Assisted Tasks
Overall, our analysis identified 32 distinct task categories in which developers use GenAI tools in their workflows. Table II presents these categories along with their definitions and usage frequencies, while Table III shows a list of examples of self-admitted GenAI usage. Unsurprisingly, excluding PR-related activities, generation tasks dominated the landscape, with code generation being particularly prominent (105 instances). Translation followed with 50 instances, while optimization and maintenance tasks accounted for 34 and 26 instances, respectively.
As mentioned above, we distinguish between developer tasks and GenAI tasks. While Table II lists the GenAI tasks, we also want to discuss human tasks related to GenAI tasks. For example, in one commit message that we analyzed (EIII) the developer acknowledged that the code was written a bit hasty on previous release due to trust in GitHub Copilot. The developer task described in the commit message was bug fixing, while the initial task that the GenAI tool supported was code generation.
@@
Recent research has shown that using AI-generated PR descriptions reduces review time and increases PR merge rates [53]. We found that developers reused generated PR descriptions as part of their commit messages. As mentioned above, this approach was very common in one particular project, which contributed 1,008 such mentions to our sample. These generated messages are not limited to this one project—similar patterns appear in popular projects such as pytorch/pytorch and hasura/graphql-engine. They are added when developers use the PR description as the message for merge or squash commits. This practice represents a form of explicit, self-admitted GenAI usage, embedding a clear marker of AI contribution directly into the software project’s official history. Note that, although this single use case contributes a large absolute count, it represents only one entry in our taxonomy of GenAI tasks (Table II) and does not affect our broader findings. To illustrate this particular use case, we include an excerpt below (EIII). Interestingly, the linked contribution guidelines (EIII) do not discuss GenAI usage.
@@
Below, we discuss the most prevalent supported tasks besides generating PR descriptions. As expected, code generation was one of the most common GenAI-supported tasks that we observed. Some self-admitted GenAI usage for code generation was straightforward, such as the following statement that we found in the source code comment documenting a method written in C#: This function was written with Chat-GPT (EIII).
Beyond code generation, developers used GenAI tools to generate other software artifacts, including test data or documentation. Besides generation, GenAI tools were also used to automate code review, for example, as part of GitHub Actions workflows (EIII): Recently, the team has added ChatGPT to GitHub Actions to perform GenAI-based code review. The specific job can be found in the chatgpt-cr.yml file.
After generation, translation emerged as the second most prevalent task in our analysis. Most mentions referred to translation between natural languages, one mention referred to translation between programming languages. An important use case was internationalization, helping developers overcome language barriers (EIII): Due to my limited proficiency in English, all English document translations are currently provided by ChatGPT, including this sentence. The one mention related code translation documented the translation of existing Python code to R (EIII): The following R code was generated using ChatGPT based on the Python code. However, the developer at the same time asked others to support them in improving the code: If anyone can refactor it to something more readable or more sensible code, please let me know or submit as a pull request.
Code optimization represented the third largest category. Developers not only acknowledged GenAI tools usage but sometimes even thanked the tools in their commit messages (EIII): Forgot tabs. Thanks, Copilot. In addition to code, GenAI tools were also used to improve UI elements (EIII): I asked chatGPT to help me brainstorm improvements to some of the labels and hint text based on the […] Interface Guidelines. I then edited them as human to improve them further. Interestingly, also in this case, the developer asked other members to review the generated content: Review and let me know if you think any are worse or weird. This, together with the human corrective actions triggered by GenAI actions we observed, points to the importance of human oversight in GenAI-assisted software development.
@@ 3.2.2 Generated Content Types
Project metadata Commit messages Target commit messages. 1,003
Source files Whole methods Target source code files, ranging from entire functions within a file. 47
Blocks within one source code file Target source code files, spanning multiple blocks within a single source code file. 45
One block within one source code file Target source code files, spanning one block within a single source code file. 39
Blocks within multiple source code files Target source code files, spanning multiple source code files. 21
Whole files Target source code files, ranging across the entire file. 12
Whole classes Target source code files, ranging across the entire class in the file. 12
Project assets Documentation files Target documentation files, which include technical documents in software projects. 106
Configuration files Target configuration files, which define the operational parameters and settings. 24
Resource files Target resource files, e.g., images, localization strings, and other binary data. 5
@@
Category Code Definition #
@@
Removal of Copilot comment Deletion of comments initially suggested by GenAI tools that are no longer relevant or correct. 2
@@
Warning Issuing cautions in the code, e.g., vulnerabilities, deprecated methods, or unstable features. 10
@@
GenAI Limitations Blame Copilot Specifically attributing errors or suboptimal code to suggestions made by a GenAI tool. 3
@@
Our analysis identified three main categories of AI-generated content in open-source software projects, organizing ten distinct codes (see Table IV). Although, as mentioned before, commit messages related to Copilot PR activities dominated our dataset with over 1,000 mentions from a single repository, examining the remaining data revealed important patterns. Developers frequently use GenAI tools to modify source files (176 mentions). However, other file types, such as documentation and configuration files, were also targeted (135 mentions).
@@
For project assets other than source code, e.g., GenAI was used to generate changelogs (EIII): Note: This changelog is improved by OpenAI ChatGPT from my broken English input. Another use case we observed was adding comments explaining options in a configuration file (EIII): These strings were provided by GitHub Copilot. I checked the first few, and they were correct.
@@ 3.2.3 Purposes of GenAI Usage
Summary RQ1: For the 1,292 GenAI mentions we analyzed, developers mainly used GenAI tools for code generation, natural language translation, and code refactoring. Source code and documentation files were the dominant generation targets. Acknowledgment of GenAI usage was a common purpose, sometimes combined with warnings about possible negative implications. Another important purpose was regulation (see RQ2). Our analysis revealed patterns of corrective actions following code generation. Our findings show that GenAI tools are actively used in open-source software and that developers are working on guiding their usage.
@@ 4 Existing Guidelines for GenAI Usage
One topic that emerged while answering RQ1 is that some open-source projects have specific policies and guidelines around GenAI usage. Therefore, as part of RQ2, we investigated how projects prohibit, restrict, or support the usage of GenAI tools. In addition to analyzing the policies and guidelines, we conducted a survey with open-source developers to understand their views on GenAI regulation.
@@ 4.1 Method
Using our sample of GenAI mentions, we found 28 mentions related to policies and usage guidelines around GenAI tool usage. We grouped them into three groups: (1) prohibitive, (2) restrictive, and (3) supportive usage. Table VI presents detailed examples drawn from 13 documentation files and commit messages in 12 GitHub repositories, where the last column indicates the number of mentions identified in the software artifact.
First, we closely examined these policies and usage guidelines to understand how exactly projects regulate GenAI usage. Second, we conducted a developer survey that included excerpts from the policies and guidelines we found. The primary goals of the survey were to: (1) collect developer perceptions on the need for GenAI tool guidance (e.g., documenting prompts or annotating generated content) and understand the actions taken on this content before integration or publication; and (2) investigate the rationale behind policies and usage guidelines. To investigate the second part, we asked participants if they contributed to one of the repositories from which we extracted policies and guidelines (see Table VI) and then showed the corresponding guidelines, asking them to elaborate on the rationale behind them. In this way, we received feedback on PVI and PVI. For developers who did not identify as contributors to one of the repositories, we showed them PVI, PVI, and PVI, asking for their feedback on those guidelines.
Our target population was the contributors of the 12 GitHub repositories in our dataset that contained explicit GenAI usage policies (see Table VI). Of these, seven had the GitHub Discussions feature enabled, which we used as our primary outreach channel to gather direct developer feedback. For the remaining five repositories where this feature was not available, as well as two repositories where our discussion posts received no response, we identified contributors on GitHub and then, to comply with GitHub’s terms of service, looked for contact details outside of GitHub (e.g., personal websites or social media profiles). We were able to determine the email addresses of 30 contributors. In total, we received eight survey responses, which we analyzed using a combination of open coding and card sorting. Informed consent was obtained. The survey questionnaire is available in our replication package [52].
@@ 4.2 Results
Table VI: Policies and guidelines for GenAI usage in software projects (RQ2): prohibitive, restrictive, and supportive policies (P); the last column (#) shows the number of mentions in the corresponding documentation file or commit message.
Goal ID Repository Excerpt #
pro P1 jqwik-team/
@@
P8 owasp/
wrongsecrets
P9 sitespeedio/
sitespeed.io
P10 theokanning/
openai-java
sup P11 avaloniaui/
@@
P13 spring-projects/
@@ 4.2.1 Policies and Usage Guidelines of GenAI Tools
As mentioned above, Table VI lists 13 software artifacts from 12 GitHub repositories that presented policies or usage guidelines for GenAI usage. We classified them into prohibitive, restrictive, and supportive policies.
Prohibitive: Policies PVI–PVI illustrate community decisions that exclude GenAI usage in the projects. Maintainers of jqwik-team/jqwik raised concerns related to the copyright situation around GenAI-generated content (PVI and PVI). Similarly, maintainers of shoelace-style/shoelace addressed ethical and licensing issues arising from the inclusion of GenAI-generated code (PVI). Regarding code reviews, maintainers of katsutedev/mal4j (PVI) and shred/acme4j (PVI) explicitly stated that contributions generated by GenAI are not acceptable. The project turms-im/turms (PVI) discouraged the use of GenAI-generated responses in discussions, citing concerns over the lack of critical thinking and responsibility. In addition, the maintainers proposed to incorporate indicators for identifying possible GenAI usage and suggested tool support, for example, a ChatGPT detector published on HuggingFace [14]. These regulations demonstrate how open-source communities are beginning to establish boundaries and safeguards to ensure responsible integration of GenAI tools within collaborative open-source software development environments.
Restrictive: Policies PVI–PVI restrict the use of GenAI tools in software development workflows without completely banning it. For example, the maintainers of graycoreio/daffodil (PVI) require developers to disclose any use of GenAI as a prerequisite for submitting a pull request. Meanwhile, maintainers of sitespeedio/sitespeed.io, theokanning/openai- java, and owasp/wrong-secrets (PVI, PVI, and PVI) advised caution when using GenAI, warning of potential inaccuracies and security risks, such as inadvertent secret leakage due to the fact that tool vendors use prompts for reinforcement learning.
Supportive: Policies PVI–PVI outline supportive guidelines for the use of GenAI tools. The maintainers of avaloniaui/avalonia (PVI) encouraged the use of GenAI to help draft pull request descriptions to facilitate the code review process. In hardisgroupcom/sfdx-hardis (PVI), maintainers recommended using GenAI for Q&A support, particularly for troubleshooting deployment issues. The project spring-projects/spring-cli (PVI) promoted the use of GenAI to generate README.md files and has even developed GenAI tooling to support automated documentation rewriting.
Overall, these policies and usage guidelines reflect a growing awareness of both the opportunities and risks of GenAI tools in open-source software projects and the willingness of the maintainers to guide their usage. An interesting direction for future work is extending the analysis to cover more policies and guidelines and correlating them with other project characteristics. For example, one could hypothesize that GPT-licensed projects are more likely to have restrictive regulations.
@@ 4.2.2 Developer Survey on GenAI Governance
Based on the analysis of the before-mentioned policies and usage guidelines, we designed ten questions regarding (i) the necessity of GenAI tool guidance; (ii) the necessity of documenting prompts and their generated contents; (iii) actions on generated content before integrating; and (iv) the rationale behind policies and usage guidelines of real-world GenAI tools. In the following, we will use $D$ to refer to individual developers who participated in our survey.
@@
Documenting Prompts and Generated Content: $D_{5,6,8}$ emphasize the importance of documenting prompts and generated content to ensure accountability in software projects. They suggest two methods to achieve this: (1) associating prompts with their functionality and sharing them under a CC BY 4.0 license, and (2) embedding prompts as code comments or in project documentation, supplemented by shared conversations, e.g., via ChatGPT links. Despite some developers considering prompt documentation unnecessary, the majority agreed that it is valuable to understand the extent of GenAI’s contributions to a project. This documentation is essential for assessing which code is potentially affected by copyright and licensing issues; it might also prove useful for later maintenance activities.
@@
Actions on Generated Content Before Integration: Developers are, compared to manually written code, more likely to perform code reviews and license compliance checks on AI-generated content before integrating it into their projects. Three developers highlighted these practices as crucial steps to ensure the quality and compliance of GenAI-based contributions. Additionally, some developers indicated that they rely on automated tools, e.g., code quality checks or automated testing, to evaluate generated content. One developer noted the importance of adding comments to the document generation context. Interestingly, $D_{2}$ explicitly stated that no additional actions are necessary, explaining that all content in the PR will be subjected to rigorous review and testing regardless. This response reflects the perspective that standard testing and code review are sufficient to ensure the quality of both AI-generated and manually created content.
@@
Project-specific GenAI Usage Guideance: The feedback we received from open-source developers regarding GenAI tool guidance reflects a combination of ethical, legal, and practical considerations. For example, the project owner of jqwik-team/jqwik ($D_{3}$) described their decision to disallow the use of GenAI tools as an ethical decision due to all its collateral damages. This statement suggests a strong position against the potential implications of accepting AI-generated contributions, with a particular focus on copyright and ethics. The regulation in the accompanying contributor agreement (PVI) explicitly prohibits contributions created using GenAI tools, citing the unresolved legal implications of training AI models on public code repositories. Similarly, the project owner of owasp/wrongsecrets ($D_{2}$) focuses on the ethical risks of using GenAI when describing the rationale behind guideline PVI. They highlight the importance of vigilance when handling sensitive data, particularly in the wrongsecrets project. They reported: This is a recommendation meant for people using WrongSecrets, and it applies more broadly than WrongSecrets or even OWASP itself. You should be conscious about what data you share, and be vigilant that you don’t input sensitive data, since tenant boundaries are mirky at best. This raises a broader concern about how user input may be stored or reused by GenAI systems. The associated recommendation emphasizes that GenAI tools often rely on reinforcement learning, which could expose sensitive data to unintended parties.
@@
Guideline PVI, which requires the disclosure of AI usage in pull requests, received support from three developers ($D_{4}$, $D_{5}$, $D_{6}$). $D_{4}$ emphasized that disclosure depends on whether a key idea was generated by AI, while $D_{5}$ highlighted the importance of transparency for license compliance. $D_{6}$ added that disclosing the percentage of GenAI involvement in contributions could reduce the likelihood of generated noise PRs and improve code review efficiency. This reflects a growing recognition of the need for transparency in collaborative software development, where understanding the role of AI in contributions can improve accountability and ensure compliance.
@@
Opinions diverge considerably for PVI, which prohibits AI-generated code. $D_{4}$ opposed such restrictions, viewing them as unnecessary limitations that could stifle productivity and innovation. $D_{6}$ criticized the policy as being overly cautious, suggesting that asking contributors to disclose percentage of generated content is sufficient. $D_{7}$ supported the regulation, noting its alignment with their own concerns about the ethical and legal implications of using GenAI tools. $D_{5}$, pointed to ethical and legal ambiguities related to AI-generated code, describing them as maintainer’s main concerns. They specifically highlighted that AI models are likely trained on large datasets that include open-source codebases with various licensing terms. $D_{4}$ and $D_{6}$’s feedback on PVI, which regulates the use of AI in community discussions, emphasizes concerns about the high false positive rates of AI identification tools and warns against deferring critical decisions to automation. $D_{5}$ argued that while LLMs are suitable for repetitive tasks and generic translations, they lack the creativity needed for meaningful contributions. This aligns with the cautionary tone of the regulation, which warns about overreliance on AI-generated content.